# Linear Discriminant Regularized Regression

### v2

@@ Abstract
Linear Discriminant Analysis (LDA) is an important classification approach. Its simple linear form makes it easy to interpret and it is capable to handle multi-class responses. It is closely related to other classical multivariate statistical techniques, such as Fisher’s discriminant analysis, canonical correlation analysis and linear regression. In this paper we strengthen its connection to multivariate response regression by characterizing the explicit relationship between the discriminant directions and the regression coefficient matrix. This key characterization leads to a new regression-based multi-class classification procedure that is flexible enough to deploy any existing structured, regularized, and even non-parametric, regression methods. Moreover, our new formulation is amenable to analysis: we establish a general strategy of analyzing the excess misclassification risk of the proposed classifier for all aforementioned regression techniques. As applications, we provide complete theoretical guarantees for using the widely used $\ell_{1}$-regularization as well as for using the reduced-rank regression, neither of which has yet been fully analyzed in the LDA context. Our theoretical findings are corroborated by extensive simulation studies and real data analysis.
Keywords: Multi-class classification, discriminant analysis, high-dimensional data, regularization, multivariate response regression, dimension reduction.
@@ 1 Introduction
This paper introduces a new regression-based approach to multi-class, high-dimensional linear discriminant analysis (LDA). The main premise that leads to LDA is that we have a sample of $n$ i.i.d. copies $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ of $(X,Y)$, with the label $Y\in\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ taking values in $L\geq 2$ classes and the conditional distributions of the feature vector $X\in\mathbb{R}^{p}$ satisfy
$X\mid Y={\bm{e}}_{\ell}~{}\sim~{}\mathcal{N}_{p}\left(\mu_{\ell},\Sigma_{w}\right),\qquad\text{for all }\ell\in[L]:=\{1,\ldots,L\}.$ (1.1)
Here $\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$ are the canonical basis vectors in $\mathbb{R}^{L}$ and the within-class covariance matrix $\Sigma_{w}$ is assumed to be strictly positive definite. For any $i\in[n]$, we say that observation $X_{i}$ belongs to class $\ell\in[L]$ if its label $Y_{i}={\bm{e}}_{\ell}$, which one may simply write as $Y_{i}=\ell$. Our insistence on the perhaps unusual notation ${\bm{e}}_{\ell}$ will become transparent later when we formulate the estimation of the optimal hyperplane as a multivariate response regression problem. Since we can always subtract the marginal mean of $X$, we assume $\mathbb{E}(X)=0_{p}$ throughout this paper mainly for simplicity, and furthermore that the class probabilities $\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$ are strictly positive for all $\ell\in[L]$. The Bayes rule under model (1.1) is to classify a new feature $x\in\mathbb{R}^{p}$ as
$g^{*}(x):=\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.2)
@@
See, for instance, Izenman (2008). Estimation of this classifier based on $(X_{i},Y_{i})$ with $i\in[n]$ becomes challenging in (i) the high-dimensional setting $p>n$ and (ii) the multi-class setting $L>2$ that allows for $L\to\infty$ as $n\to\infty$.
@@
Few approaches consider (ii) but notable exceptions are Abramovich and Pensky (2019); Nibbering and Hastie (2022). We recall various approaches to deal with (i). We will use the following notation. Write
$\pi=(\pi_{1},\ldots,\pi_{L})^{\top},\quad D_{\pi}={\rm diag}(\pi)\in\mathbb{R}^{L\times L},\quad{M}=(\mu_{1},\ldots,\mu_{L})\in\mathbb{R}^{p\times L}$ (1.3)
and denote the unconditional covariance matrix of $X$ by
$\Sigma={\rm Cov}(X)=\Sigma_{w}+{M}D_{\pi}{M}^{\top}.$ (1.4)
We call $\Sigma_{w}$ the within-class covariance matrix and $\Sigma_{b}:=\Sigma-\Sigma_{w}={M}D_{\pi}{M}^{\top}$ the between-class covariance matrix. The latter has rank less than $L$ due to $\mathbb{E}(X)=0_{p}$.
@@
In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure by approximately solving the equation $\Sigma_{w}\beta^{*}=\mu_{2}-\mu_{1}$ for $\beta^{*}$ to estimate this direction. Other procedures based on the same or similar equations include, for instance, Qiao et al. (2009); Shao et al. (2011); Fan et al. (2012) and Tibshirani et al. (2002); Fan and Fan (2008); Cai and Zhang (2019); Chen and Sun (2022) for $L>2$. However, it isn’t always reasonable to assume that such directions are sparse and the procedure is less attractive compared to regression-based approaches discussed below shortly. Witten and Tibshirani (2011) alternatively use the fact that the Bayes rule (1.2) is equivalent with Fisher’s discriminant rule
$\operatorname*{arg\,min}_{\ell\in[L]}~{}(x-\mu_{\ell})^{\top}F{F^{\top}}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.5)
Here the columns of the matrix $F=(F_{1},\ldots,F_{K})\in\mathbb{R}^{p\times K}$, for some $K<L$, are implicitly defined via, for $1\leq k\leq K$,
@@
$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}~{}\beta^{\top}\Sigma_{b}\beta\quad\text{subject to}&~{}\beta^{\top}\Sigma_{w}\beta=1,~{}\beta^{\top}\Sigma_{w}F_{i}=0,~{}~{}\forall~{}i<k.\end{split}$ (1.6)
@@
Witten and Tibshirani (2011) develop methodology tailored to the high-dimensional scenarios with a diagonal matrix $\Sigma_{w}$ and sparse columns of $F$. However, the implicit formulation (1.6) of the matrix $F$ makes it extremely difficult to obtain the global solution and to derive theoretical properties of the regularized estimator and the resulting classifier.
@@
More closely related to our approach is optimal scoring which dates back to De Leeuw et al. (1976); Young et al. (1978) and regains attention in the LDA setting, see Hastie et al. (1995, 1994); Clemmensen et al. (2011); Gaynanova (2020) and the references therein. By writing the response matrix and feature matrix as
${\mathbf{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L},\quad{\mathbf{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p},$ (1.7)
optimal scoring is a regression-based approach that aims to solve a sequence of optimization problems: for $1\leq k\leq K<L$,
@@
Here $\widehat{\theta}_{i}$ is the solution to (1.8) in the $i$th iteration. In the low-dimensional case ($p<n$), Hastie et al. (1995) show that the solutions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$ from (1.8) can be computed from canonical correlation analysis (CCA), and are further related, via CCA, with the matrix $F$ obtained from the sample analogue of (1.6). As a result, classification can be based on the dimension reduction directions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$. In the high-dimensional setting, regularization in (1.8) is needed. For instance, Hastie et al. (1995) consider
$\begin{split}\min_{\theta\in\mathbb{R}^{L},\ \beta\in\mathbb{R}^{p}}&~{}\|{\mathbf{Y}}\theta-{\mathbf{X}}\beta\|_{2}^{2}+\beta^{\top}\Omega\beta\quad\text{s.t.}~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\theta=n,~{}\theta^{\top}{\mathbf{Y}}^{\top}{\mathbf{Y}}\widehat{\theta}_{i}=0,~{}\forall~{}i<k,\end{split}$ (1.9)
for a chosen $p\times p$ symmetric, positive definite matrix $\Omega$, and show that the global solution to (1.9) can be computed. A possible choice of $\Omega=\lambda\mathbf{I}_{p}$ for some $\lambda>0$ leads to the ridge penalty (Campbell, 1980; Friedman, 1989). To accommodate more general regularization, Gaynanova (2020) studies the matrix formulation of (1.9),
@@
and proves that the global solution to (1.10) can be computed, provided pen$(B)$ satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$. Gaynanova (2020) focuses on the group-lasso penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ and analyzes the resulting classifier. However, it is not clear how the analysis can be extended to other penalized estimation. Moreover, the requirement that pen($B$) is orthogonal invariant excludes the familiar lasso penalty $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ as well as its extensions such as the elastic net penalty. That being said, for penalties that are not orthogonal invariant, there is no guarantee of computing the global solution to (1.10), let alone any theoretical property of the resulting estimator.
Key to our approach is the following reformulation of the problem. We show in Lemma 1 of Section 2 that the discriminant directional matrix under model (1.1)
@@
not only share the same column space, but have a closed-form connection, to wit,
@@
for some invertible $L\times L$ matrix $H$ with explicit expression. This connection is novel and distinguishes our approach from the current literature. Although the fact that columns of $B$ and $B^{*}$ share the same span is known (see, for instance, Hastie et al. (1995); Ye (2007); Lee and Kim (2015); Nie et al. (2022) and the references therein), their closed-form connection given by (1.13) has not been discovered, and a thorough theoretical investigation on the corresponding classifier has not been conducted. To the best of our knowledge, only in the binary case ($L=2$), Mai et al. (2012) leverages sparse linear regression between an appropriately encoded label vector and the feature matrix ${\mathbf{X}}$ for classification and proves consistency. Extension of such procedure to handle multi-class responses with high-dimensional features has been a longstanding open problem. In view of (1.13), we propose to estimate the discriminant direction matrix $B^{*}$ by first estimating the regression matrix $B$ as
$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{}{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B),$ (1.14)
and then estimating $H^{-1}$ based on $\widehat{B}$ via its explicit expression in Lemma 1. The advantages of our approach as well as our main contributions are three-fold:
@@
    Akin to optimal scoring, our approach in (1.14) is regression-based, which, as pointed out in Hastie et al. (1995, 1994), is often easier to compute and more amenable to extensions to regularized and non-parametric regression as well as model selection than canonical correlation analysis, Fisher’s discriminant rule and the aforementioned procedures based on approximately solving the equation $\Sigma_{w}B^{*}=M$ for $B^{*}$. We can also easily tap into the large literature on estimation of the high-dimensional regression matrix $B\in\mathbb{R}^{p\times L}$ in (1.14).
@@
    After computing $\widehat{B}$ from (1.14), we only need to estimate the inverse of an $L\times L$ matrix to estimate the discriminant direction matrix $B^{*}$ rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context. Our approach in Section 2.2 uses the explicit form of $H$ in Lemma 1 and enjoys computational advantage as well as numerical stability, compared to the existing approaches. In Section 2.3, we further offer an alternative way to estimate the Bayes rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$. This leads to a practically useful method, which could extract fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization.
  3. (c)
@@
    We are able to quantify the errors of estimating both $B$ and $B^{*}$ as well as to provide theoretical guarantees for the resulting classifier, thanks to (1.13). Concretely, in Section 3.1 and Section 3.2 we provide a general strategy for analyzing the excess misclassification risk of the proposed classifier, which is valid for a generic estimator $\widehat{B}$ of $B$. While various choices of $\widehat{B}$ are feasible in our approach, more specifically in (1.14), we apply the general result to two popular estimators: the lasso estimator in Section 3.3 and the reduced rank estimator in Section 3.4. The former also contrasts our approach with Cai and Zhang (2019) and Witten and Tibshirani (2011) which assume sparsity of $B^{*}$ and $F$, respectively. The latter is particularly suitable when the rank of the discriminant direction matrix $B^{*}$ is low and the number of classes $L$ is moderate / large. For instance, our earlier work Bing and Wegkamp (2023) considers a multi-class classification setting with high-dimensional features $X\in\mathbb{R}^{p}$ that follow a factor model $X=AZ+W$ with unobserved (latent) features $Z\in\mathbb{R}^{r}$ with $r<p$ and random noise $W$, that is independent of both $Z$ and the label $Y$. In this case, the discriminant direction matrix $B^{*}$ has rank no greater than $r$ when $r<L$. The procedure in Bing and Wegkamp (2023), however, is different than the one proposed here. We refer to Remark 1 in Section 2.2 for more discussion.
@@
Finally, we provide a thorough simulation study in Section 4 to corroborate our theoretical findings. In Section 5, we report our findings in several real data studies. The Appendix contains most of the proofs and additional simulation study.
Notation.

### v3

@@ Abstract
Linear Discriminant Analysis (LDA) is an important classification approach. Its simple linear form makes it easy to interpret, and it is well-suited for handling multi-class responses. It is closely related to other classical multivariate statistical techniques, such as Fisher’s discriminant analysis, canonical correlation analysis and linear regression. In this paper we strengthen its connection to multivariate response regression by characterizing the explicit relationship between discriminant directions and regression coefficients. This key characterization leads to a new regression-based multi-class classification procedure that is flexible enough to deploy any existing structured, regularized, and even non-parametric, regression methods. In contrast with the existing regression-based approaches, our new formulation is particularly amenable to establish provable guarantees: we establish a general strategy of analyzing the excess misclassification risk of the proposed classifier for all aforementioned regression techniques. As applications, we provide complete theoretical guarantees for using the widely used $\ell_{1}$-regularization as well as for using the reduced-rank regression, neither of which has yet been fully analyzed in the LDA context. Our theoretical findings are corroborated by extensive simulation studies and real data analysis.
Keywords: Dimension reduction, discriminant analysis, high-dimensional data, multi-class classification, multivariate response regression, regularization.
@@ 1 Introduction
Linear Discriminant Analysis (LDA) is a popular tool for predicting a categorical response using a set of explanatory variables. Its popularity stems from several favorable properties, such as ease of interpretation, reasonable robustness to departures from normality, and the capability to handle responses with multiple classes. See, Hastie et al. (1995) and many references in Seber (2009). This paper introduces a new regression-based approach to multi-class, high-dimensional linear discriminant analysis (LDA).
Assume that $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ are i.i.d. copies of a random pair $(X,Y)$, where the label $Y$ takes values in $\{{\bm{e}}_{1},\ldots,{\bm{e}}_{L}\}$, the canonical basis vectors in $\mathbb{R}^{L}$, with $L\geq 2$ classes, and the feature vector $X\in\mathbb{R}^{p}$ has conditional means collected in the $p$ by $L$ matrix ${M}=(\mu_{1},\ldots,\mu_{L})$ and the within-class covariance matrix
$\Sigma_{w}={\rm Cov}(X\mid Y={\bm{e}}_{\ell}),\quad\text{for all }\ell\in[L]:=\{1,\ldots,L\}.$
We assume $\Sigma_{w}$ to be strictly positive definite and the class probabilities $\pi_{\ell}=\mathbb{P}\{Y={\bm{e}}_{\ell}\}$, for all $\ell\in[L]$, are strictly positive. We denote by $\Sigma$ the unconditional covariance matrix of $X$. Since we can always subtract the marginal mean of $X$, we assume $\mathbb{E}(X)=0_{p}$.
For a new feature $x\in\mathbb{R}^{p}$, LDA predicts its corresponding label as
$\operatorname*{arg\,min}_{\ell\in[L]}~(x-\mu_{\ell})^{\top}\Sigma_{w}^{-1}(x-\mu_{\ell})-2\log(\pi_{\ell}).$ (1.1)
@@
The LDA rule in (1.1) coincides with the Bayes rule when $X\mid Y$ is Gaussian. See, for instance, Izenman (2008). Estimation of the classifier (1.1) based on $(X_{1},Y_{1}),\ldots,(X_{n},Y_{n})$ becomes challenging in (i) the high-dimensional setting $p>n$ and (ii) the multi-class setting $L>2$ that allows for $L\to\infty$ as $n\to\infty$.
1.1 Existing approaches
@@
Few approaches consider (ii) but notable exceptions are Levy and Abramovich (2023); Abramovich and Pensky (2019); Nibbering and Hastie (2022).
@@
We recall various approaches to deal with (i). In high-dimensional settings $p>n$, Cai and Liu (2011) assume that the direction $\beta^{*}:=\Sigma_{w}^{-1}(\mu_{2}-\mu_{1})$ is sparse in the binary case $L=2$, and propose a Dantzig-type procedure to estimate this direction by approximately solving the equation $\Sigma_{w}\beta^{*}=\mu_{2}-\mu_{1}$ for $\beta^{*}$. Other procedures based on the same or similar equations include, for instance, Qiao et al. (2009); Shao et al. (2011); Fan et al. (2012) for $L=2$, and Tibshirani et al. (2002); Fan and Fan (2008); Cai and Zhang (2019); Chen and Sun (2022); Mai et al. (2019); Gaynanova et al. (2016) for $L>2$. Specifically, for multi-class responses, there are L1L-1 directions $\beta_{\ell}^{*}=\Sigma_{w}^{-1}(\mu_{\ell}-\mu_{1})$ for $2\leq\ell\leq L$, which can be solved via
$(\beta_{2}^{*},\ldots,\beta_{L}^{*})=\operatorname*{arg\,min}_{\beta_{2},\ldots,\beta_{L}}~\sum_{\ell=2}^{L}\left({1\over 2}\beta_{\ell}^{\top}\Sigma_{w}\beta_{\ell}-\beta_{\ell}^{\top}(\mu_{\ell}-\mu_{1})\right).$ (1.2)
For large $p>n$, sparsity is imposed on $\beta_{2}^{*},\ldots,\beta_{L}^{*}$. These directions can be estimated by replacing $\Sigma_{w}$ and $\mu_{\ell}$ in (1.2) by their sample counterparts, and adding a penalty that encourages sparsity (Mai et al., 2019; Gaynanova et al., 2016; Wang et al., 2021; Zeng et al., 2024). However, it is not always reasonable to assume that the directions $(\beta_{2}^{*},\ldots,\beta_{L}^{*})$ are sparse, especially when $\Sigma_{w}$ has many non-negligible off-diagonal entries. Moreover, the above procedure–based on solving a quadratic program–is less appealing than the regression-based approaches discussed below in (1.5), (1.6) and (1.10), particularly in high-dimensional settings where structural or penalized estimation is desired and issues such as tuning parameter selection and computational efficiency become more critical. Also, see part (a) of our main contributions in Section 1.2.
Witten and Tibshirani (2011) alternatively use the fact that the first term in the LDA rule (1.1) is equivalent with Fisher’s discriminant rule
$\operatorname*{arg\,min}_{\ell\in[L]}~(x-\mu_{\ell})^{\top}F{F^{\top}}(x-\mu_{\ell}),$ (1.3)
where the columns of $F=(F_{1},\ldots,F_{K})\in\mathbb{R}^{p\times K}$, for some $K<L$, are defined via, for $k\in[K]$,
@@
$\begin{split}F_{k}:=\operatorname*{arg\,max}_{\beta\in\mathbb{R}^{p}}~\beta^{\top}\Sigma\beta\quad\text{subject to}&~\beta^{\top}\Sigma_{w}\beta=1,~\beta^{\top}\Sigma_{w}F_{i}=0,~~\forall~i<k.\end{split}$ (1.4)
@@
Criteria similar to (1.3) – (1.4) and their connection to discriminant analysis have been considered in Safo and Ahn (2016); Ahn et al. (2021); Jung et al. (2019). Witten and Tibshirani (2011) develop methodology tailored to the high-dimensional scenarios with a diagonal matrix $\Sigma_{w}$ and sparse columns of $F$. However, the formulation (1.4) of the matrix $F$ makes it extremely difficult to obtain the global solution and to derive theoretical properties of the regularized estimator and the resulting classifier.
@@
More closely related to our approach is optimal scoring which dates back to De Leeuw et al. (1976); Young et al. (1978) and regains attention in the LDA setting, see Hastie et al. (1995, 1994); Clemmensen et al. (2011); Gaynanova (2020) and the references therein. By writing the response matrix and feature matrix as ${\mathbf{Y}}=(Y_{1},\ldots,Y_{n})^{\top}\in\{0,1\}^{n\times L}$ and ${\mathbf{X}}=(X_{1},\ldots,X_{n})^{\top}\in\mathbb{R}^{n\times p}$, optimal scoring is a regression-based approach that aims to solve a sequence of optimization problems: for $1\leq k\leq K<L$,
@@
Here $\widehat{\theta}_{i}$ is the solution to (1.5) in the $i$th iteration. In the low-dimensional case ($p<n$), Hastie et al. (1995) show that the solutions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$ from (1.5) can be computed from canonical correlation analysis (CCA), and are further related, via CCA, with the matrix $F$ obtained from the sample analogue of (1.4). As a result, classification can be based on the dimension reduction directions $\widehat{\beta}_{1},\ldots,\widehat{\beta}_{K}$. In the high-dimensional setting, regularization is needed. For instance, Hastie et al. (1995) considers adding $\beta^{\top}\Omega\beta$ to the loss function of (1.5) for a chosen $p\times p$ symmetric, positive definite matrix $\Omega$, and show that the global solution can be computed. A possible choice of $\Omega=\lambda\mathbf{I}_{p}$ for some $\lambda>0$ leads to the ridge penalty (Campbell, 1980; Friedman, 1989). To accommodate more general regularization, Gaynanova (2020) studies the matrix formulation of (1.5),
@@
and proves that the global solution to (1.6) can be computed, provided pen(B)(B) satisfies $\text{pen}(BR)=\text{pen}(B)$ for any orthogonal $K\times K$ matrix $R$. Gaynanova (2020) focuses on the group-lasso penalty $\|B\|_{1,2}=\sum_{j=1}^{p}\|B_{j\cdot}\|_{2}$ (which is orthogonal invariant) and analyzes the resulting classifier when $X\mid Y$ is Gaussian. However, it is not clear how the analysis can be extended to other penalized estimation. In particular, the requirement that pen($B$) is orthogonal invariant excludes the familiar lasso penalty $\|B\|_{1}=\sum_{j,k}|B_{jk}|$ and the elastic net penalty. Worse, for penalties that are not orthogonal invariant, there is no guarantee of computing the global solution to (1.6), let alone any theoretical property of the resulting estimator.
1.2 Our contributions
We propose a new regression based multi-class classification approach that is suitable in both low- and high-dimensional settings and has provable theoretical guarantees. Key to our approach is the following reformulation of the problem. We show in Lemma 1 of Section 2 that the discriminant directional matrix in (1.1)
@@ 1.2 Our contributions
not only share the same column space, but have a closed-form connection, specifically,
@@
for some invertible $L\times L$ matrix $H$, which we derive explicitly. This connection is novel and distinguishes our approach from the current literature. While existing literature has leveraged the fact that $B$ and $B^{*}$ share the same column space to develop two-step classification procedures, typically by first estimating this shared subspace (see, e.g., Hastie et al. (1995); Ye (2007); Lee and Kim (2015); Nie et al. (2022); Gaynanova et al. (2016); Ahn et al. (2021)), these approaches lack theoretical guarantees for the resulting classifier. In contrast, our procedure is built upon the new closed-form characterization of $B$ given in (1.9), which not only leads to a straightforward classifier but also enables a rigorous theoretical analysis of its misclassification risk. To the best of our knowledge, only in the binary case ($L=2$), Mai et al. (2012) employs sparse linear regression between an appropriately encoded label vector and the feature matrix $\mathbf{X}$} for classification and proves consistency of their classifier. Extension of such regression formulation to handle multi-class responses with high-dimensional features has been a longstanding open problem.
In view of (1.9), we propose to estimate the discriminant direction matrix $B^{*}$ by first estimating the regression matrix $B$ as
$\widehat{B}=\operatorname*{arg\,min}_{B\in\mathbb{R}^{p\times L}}~{1\over n}\|{\mathbf{Y}}-{\mathbf{X}}B\|_{F}^{2}+\text{pen}(B)$ (1.10)
and then estimating $H^{-1}$ based on $\widehat{B}$ via its explicit expression in Lemma 1. The advantages of our approach as well as our main contributions are two-fold:
@@
    From a methodological perspective, our approach in (1.10), akin to optimal scoring, is regression-based. As pointed out by Hastie et al. (1995, 1994), regression-based methods are generally much easier to compute – especially in high-dimensional settings – than canonical correlation analysis (Gaynanova, 2020), Fisher’s discriminant rule (Witten and Tibshirani, 2011), and the aforementioned procedures based on approximately solving in (1.2) as in (Mai et al., 2019; Cai and Zhang, 2019). Moreover, regression-based approaches are more amenable to extensions involving regularized and non-parametric regression, as well as model selection. We can also easily tap into the large literature on estimation of the high-dimensional regression matrix $B\in\mathbb{R}^{p\times L}$ in (1.10). However, we caution that the nonlinear relation $\mathbb{E}[{\mathbf{Y}}\mid{\mathbf{X}}]$ and $\mathbf{X}$} complicates the theoretical analysis.
    After computing $\widehat{B}$ from (1.10), we only need to estimate the inverse of an $L\times L$ matrix to estimate the discriminant direction matrix $B^{*}$ rather than that of a $p\times p$ matrix $\Sigma_{w}$ as in the classical LDA context. Our approach in Section 2.2 uses the explicit form of $H$ in Lemma 1 and enjoys both computational advantage and numerical stability, compared to the existing approaches. In Section 4.1, we further offer an alternative way to estimate the LDA rule that performs Fisher’s discriminant analysis on low-dimensional transformations $\widehat{B}^{\top}x\in\mathbb{R}^{L}$ of the original features $x\in\mathbb{R}^{p}$. This leads to a practically useful method, which could extract fewer discriminant directions than those in $B^{*}$, for both downstream analysis and visualization (see, for instance, the real data analysis in Section 5).
@@
    Theoretically, unlike the intractable analysis of optimal scoring with general regularization, we are able to quantify the errors in estimating both $B$ and $B^{*}$, and provide theoretical guarantees for the resulting classifier, thanks to (1.9). Concretely, under the assumption that the distributions of $X\mid Y={\bm{e}}_{\ell}$ are Gaussian $\mathcal{N}_{p}(\mu_{\ell},\Sigma_{w})$, we provide in Sections 3.1 and 3.2 a general strategy for analyzing the excess misclassification risk of the proposed classifier, which is valid for any generic estimator of $B^{*}$, including all existing estimators as well as our proposed estimator $\widehat{B}^{*}=\widehat{B}\widehat{H}^{-1}$. While various choices of $\widehat{B}$ are feasible in our approach, more specifically in (1.10), we apply the general result to two popular estimators: the lasso estimator in Section 3.3 and the reduced rank estimator in Section 3.4. The former also contrasts our approach with Cai and Zhang (2019) and Witten and Tibshirani (2011) which assume sparsity of $B^{*}$ and $F$, respectively. The latter is particularly suitable when the rank of the discriminant direction matrix $B^{*}$ is low and the number of classes $L$ is moderate / large. For instance, our earlier work Bing and Wegkamp (2023) considers a multi-class classification setting with high-dimensional features $X\in\mathbb{R}^{p}$ that follow a factor model $X=AZ+W$ with unobserved (latent) features $Z\in\mathbb{R}^{r}$ with $r<p$ and random noise $W$, that is independent of both $Z$ and the label $Y$. In this case, the discriminant direction matrix $B^{*}$ has rank no greater than $r$ when $r<L$. The procedure in Bing and Wegkamp (2023), however, is different than the one proposed here. We refer to Remark 1 in Section 2.2 for more discussion.
@@
Finally, we provide a thorough simulation study in Section 4 to corroborate our theoretical findings. Section 5 contains our findings in several real data studies. The Supplement (Bing et al., 2025) contains all the proofs.
1.3 Notation